OpenStudy (anonymous):

280 divided by 40 =??? with remainder if there is one

OpenStudy (anonymous):

70

OpenStudy (anonymous):

there will b zero remainder

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

7

OpenStudy (anonymous):

7

OpenStudy (anonymous):

u for got a 0

OpenStudy (anonymous):

142343 sometimes over confidence is detrimental !!!!!

OpenStudy (anonymous):

40*70 = ?

OpenStudy (anonymous):

280

OpenStudy (anonymous):

duh

OpenStudy (anonymous):

40=10*4 70=10*7 40*70=10*4*10*7= 4*7 * 10*10 = 28*100 = 2800

OpenStudy (anonymous):

quotient is 7, while remainder is 0